Rick's bottle of cough syrup is 3/4 full. He gives 1/6 of the medication remaining in the bottle to his daughter. What fraction of the whole bottle of cough syrup did Rick give his daughter?
Answer by TimothyLamb(4379)

---
fraction of whole:
(3/4)(1/6) = (3/24) = 1/8
---
answer: 1/8
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php